# High-Fidelity And Complex Test Data Generation For Real-World SQL Code Generation Services

### v2

@@ Abstract
The demand for high-fidelity test data is paramount in industrial settings where access to production data is largely restricted. Traditional data generation methods often fall short, struggling with low-fidelity and the ability to model complex data structures and semantic relationships that are critical for testing complex SQL code generation services like Natural Language to SQL (NL2SQL). In this paper, we address the critical need for generating syntactically correct and semantically “meaningful” mock data for complex schema that includes columns with nested structures that we frequently encounter in Google SQL code generation workloads. We highlight the limitations of existing approaches used in production, particularly their inability to handle large and complex schema, as well as the lack of semantically coherent test data that lead to limited test coverage. We demonstrate that by leveraging Large Language Models (LLMs) and incorporating strategic pre- and post-processing steps, we can generate realistic high-fidelity test data that adheres to complex structural constraints and maintains semantic integrity to the test targets (SQL queries/functions). This approach supports comprehensive testing of complex SQL queries involving joins, aggregations, and even deeply nested subqueries, ensuring robust evaluation of SQL code generation services, like NL2SQL and SQL Code Assistant services. Our results demonstrate the practical utility of an out-of-the-box LLM (gemini) based test data generation for industrial SQL code generation services where generating realistic test data is essential due to the frequent unavailability of production datasets.
@@ I Introduction
One key challenge for effective test data generation lies in achieving realism and diversity within industrial contexts. Mock data must be sufficiently realistic and diverse to effectively simulate the wide range of scenarios encountered in real-world production systems, including both common and rare edge cases. Existing data generation methods often fall short in capturing this diversity, while rule-based systems tend to be inflexible and limited in their ability to generate varied data that reflects the complexities of industrial data distributions [9, 10]. Another challenge is supporting test cases with complex data structure, consistency, and data integrity requirements [5, 4, 11, 3]. For industrial applications heavily reliant on structured data types such as JSON, XML, protocol buffers (a.k.a. proto [12]), SQL and Open API schemas, it is paramount that the generated mock data strictly adheres to the specified structure and maintains data integrity and consistency [13, 14]. Specifically, our SQL code generation services utlizie GoogleSQL [15] that supports complex nested data types, like array, struct and proto – making it a representative case for the challenges faced by many modern SQL dialects that support nested data structures. Manual processes are highly susceptible to errors that can compromise the validity of testing and development. Furthermore, scalability and efficiency of generation are critical for enterprise needs, especially when a large amount of test data is needed [16, 17, 10]. As enterprise software systems grow in scale and complexity, the need for large volumes of mock data increases exponentially. Manual and heuristic data generation approaches become impractical and unsustainable in such environments, hindering development velocity and quality assurance efforts.
@@
The emergence of Large Language Models (LLMs) presents a paradigm shift, offering new avenues for automating the generation of mock data, especially for structured formats such as schemas and protocol buffers. The use of LLMs for synthetic data generation is becoming popular in various domains [18, 16, 19, 20, 21, 22, 23, 24], the generation of structured data introduces unique challenges that are particularly relevant to industrial applications. LLMs, when appropriately trained, can provide realistic and privacy-preserving data synthesis. They can generate mock data that more closely approximates real-world textual data distributions and patterns, capturing subtle nuances and semantic relationships between data attributes that are difficult to replicate with traditional methods [25, 26]. This is crucial for realistic testing while also avoiding the privacy risks associated with using actual production data (regulatory security and privacy policy restricts access to user data). Additionally, LLMs can understand SQL structures and facilitate structurally compliant and schema-aware data generation [18, 27, 26, 28]. Given a schema or proto definition, LLMs can be guided to generate mock data that adheres to the specified structure and constraints. However, tabular data generation via language models remains an active and evolving field, facing challenges such as accurately capturing complex inter-column (pairwise) relationships, mitigating the risk of “hallucinations” or data inconsistencies, and effectively handling rare values and outliers (also see Table II for vanilla LLM evaluation results) [26, 28].
Deep generative models [29, 30, 31] have shown promise in tabular data generation, particularly for categorical, numerical, and text attributes. However, they are fundamentally limited by their focus on flat schemas, neglecting nested structures, and their reliance on training data, rendering them infeasible in scenarios where production data is restricted. In contrast, our LLM-based approach, leveraging gemini [32] out-of-the-box, directly addresses these limitations by generating structurally compliant data for arbitrarily complex, nested schemas without requiring fine-tuning.
@@
However, current LLM-based data generation approaches often fall short in several critical areas. A significant limitation lies in their ability to effectively handle more complex data structures commonly found in our workloads. Specifically, LLMs struggle to accurately generate data that conforms to large (many columns or fields) and complex nested structures. While LLMs can generate data that adheres to basic schemas, accurately generating data that conforms to deeply nested structures or schemas with complex dependencies remains a challenge. This limitation is particularly problematic in testing real-world SQL generation services that rely on intricate data models, where the nuances of nested relationships over very wide tables are critical for realistic testing and validation.
@@
Furthermore, ensuring the semantic validity of the generated data is another crucial challenge that must be addressed to generate truly realistic test data. This is even more critical for testing generated SQL queries (e.g., in NL2SQL systems) on user data. While LLMs might generate data that is syntactically correct and structurally valid, it often lacks the necessary semantic coherence across columns and tables. Most test data generation approaches focus on approximating per-column or pairwise column distributions from the production data [29, 30, 21, 9, 1, 3, 4, 5] leading to illogical combinations of column values, semantically incoherent join tables – violating real-world business rules. Overcoming these limitations, including both structural and semantic challenges, is crucial for fully realizing the potential of LLMs in generating comprehensive and accurate mock data for complex workloads.
@@
In this paper, we address the critical need for generating syntactically correct and semantically “meaningful” mock data for structurally well-defined data formats, focusing on the specific requirements of industrial use cases. We define “meaningful” data as data that effectively supports a wide range of complex test cases for Google SQL code generation services, like NL2SQL and SQL code assistance systems. Robust evaluation of such generation systems requires realistic and complex test data, which are often inaccessible due to privacy or lack of complex data generation capabilities. Services, like NL2SQL, support arbitrarily complex use cases, generating intricate SQL queries involving joins, aggregations, and nested subqueries—scenarios prevalent in real-world applications. The meaningful test data generation should also include the ability to generate data that respects various constraints, such as generating data within a specific date range when the SQL query under test includes a corresponding filter (e.g., WHERE year = 2022). Moreover, we emphasize the importance of maintaining column correlations and data dependencies within a schema, which are often crucial for simulating realistic data behavior. While the ability to generate complex, multi-faceted datasets is essential for thoroughly evaluating these queries, semantic integrity of test data is also critical, in that incorrect or illogical data can lead to misleading test results, failing to expose critical flaws in the generated SQL queries. Without such robust testing, the reliability and trustworthiness of SQL code generation services are severely compromised.
@@
  • LLM-based High-Fidelity and Efficient Test Data Generation: We present a novel approach leveraging LLMs, enhanced with pre- and post-processing, to generate high-fidelity test data compliant with complex and large schemas, including nested column types, addressing the limitations of existing data generation methods in production. Our generation method is significantly more efficient than manual generation used in production where production data access is restricted.
  • Enhanced Test Coverage for Complex SQL Queries: We demonstrate that our generated test data significantly improves test coverage of Google SQL code generation services, where many test cases are disabled due to limited access to production data and the difficulties of manual data generation for large, complex schemas. Our fundamental approach is dialect-agnostic and should be readily applicable to other SQL code generation services that face similar challenges.
  • Semantic Coherence With Test SQL Queries: We emphasize and achieve the generation of semantically meaningful and validated test data. We provide a comparative analysis against production data sampling, showcasing the ability of our generated test data to provide meaningful tests that cover edge cases and semantically relevant predicate columns.
@@ II SQL code generation services
In this section, we examine the challenges of generating meaningful test data for two real-world GoogleSQL [15] SQL code generation services: SQL Code Assistant and NL2SQL. These services represent distinct use cases with varying complexities, highlighting the limitations of traditional test data generation methods which produces data that yields low coverage of the GoogleSQL queries under test.
TABLE I: Traditional test data generation methods for real-world SQL code generation services. Test coverage is largely limited by the test data quality, especially when the underlying data structure/schema is large and complex with deep nesting.
@@
Table I summarizes test & schema complexities, as well as the traditional test data generation methods used for the services. The underlying data schema for both services are complex with many columns/fields with nested fields (nested proto and array types). The test complexity for NL2SQL is even more complex as the service generates SQL queries spanning over multiple tables across many customer domains. Test cases are submitted by the internal developers and also by crowdsourcing. The test SQL queries for SQL Code Assistant are mostly simple (single-table) selection queries; however, the manually curated test data overlooks complex nested structures from the schema limiting the test cases to SQL queries targeting mostly flat schemas with non-nested fields. The SQL query complexity vary, but most of them use nested fields and subqueries with joins. SQL code generation services require high-fidelity test data that adheres to the complex schema with nested structures that support complex and diverse sets of SQL queries/functions. Additionally, NL2SQL service evaluation requires testing if predicted SQL generates a result set that is semantically coherent with the natural language question. This also means that the test data instances should be realistic and aligned with the context of the posed data question. Unfortunately, production sampling, though anonymized, is restricted to a few accessible domains and tables, resulting in incomplete test scenarios and low test coverage.
@@ II-A SQL Code Assistant
1CREATE PUBLIC FUNCTION GetBalanceInUsd(Input fake_table) 2RETURNS INT64 3AS ( 4 CASE 5 WHEN private_info.running_balance.currency = ’USD’ THEN private_info.running_balance.amount 6 WHEN private_info.running_balance.currency = ’GBP’ THEN private_info.running_balance.amount * 1.26 7 WHEN private_info.running_balance.currency = ’EUR’ THEN private_info.running_balance.amount * 1.05 8 ELSE NULL 9 END 10);
@@ II-B NL2SQL
1Question: Which fake_devices have tasks related to fake_task_name with a start date after 2020-06-01, and what are the tasks and their start dates for the fake_devices? 2---------------------------------------------------- 3 4SELECT 5 name, 6 ARRAY_AGG(STRUCT(task, start_date)) 7FROM ( 8 SELECT 9 d.name, 10 t.name AS task, 11 FORMAT_TIMESTAMP("%Y-%m-%d", PARSE_TIMESTAMP("%Y-%m-%dT%H:%M:%E*S%Ez", t.start_date)) AS start_date 12 FROM fake_project_tasks AS t 13 INNER JOIN fake_devices AS d 14 ON t.project_id = d.id 15 WHERE ( 16 t.name LIKE ’%fake_task_name’ 17 ) 18 AND t.start_date > "2020-06-01" 19 ORDER BY start_date 20) 21GROUP BY name;
@@ II-C Problem Definition
Let $\Lambda$ be a schema definition, like table schema or proto, with various types of fields/columns $\Xi$. The column set $\Xi$ includes primitive type (e.g., numerical, string, boolean) $\xi^{prim}$, categorical type $\xi^{cat}$ (e.g., enum), and nested type (e.g., struct, array, proto) $\xi^{nest}$ . $\xi^{nest}$ can be another (nested) schema definition $\Lambda_{i}$ that can have more nested structures within $\xi^{nest}_{i}$. Our work focuses on generating arbitrarily complex test data with any number of columns and with deeply nested structures, $\Xi=\{\xi^{prim}_{1},\xi^{cat}_{2},\xi^{nest}_{3},...,\xi_{n}\}$, for testing test SQL queries $\Omega=\{\omega_{1},\omega_{2},...,\omega_{m}\}$ and also without any training tabular dataset $\tau_{\Lambda}$ that confirms to $\Lambda$ – restricted production data access in real-world SQL code generation services.
@@ III-A Pre-processor
This module will seed the signals needed to generate meaningful data, automatically based on information from the context map and the test SQL query body. Additionally this can also be passed from the user and any user provided input will override auto-generated ones. The signal will be a map of column name to column value (literal value, range or values, or a generator like incremental or other distribution, e.g., use a uniform distribution from 0 to 1 to populate this numeric column). These signals would be added to the context map and used later for post-processing.
Signals for Data Generation: Additionally the following signals would be needed to generate meaningful data:
@@
[…]
Figure 4: Pre-processor parses out subqueries (middle) from the original SQL query (top) for seeding and independent and dependent data generation. Join information is extracted to highlight the table and column association for the join (bottom).
@@ III-B Context extractor
This module generates context relevant to the schema. We use a separate system to collect human written comments for a schema and proto from comments in data schema, source code and documentations. These comments are available at the schema/proto level and also individual column level. We use LLM to make these comments consistent and fill in missing columns. For protos, by themselves or as a column in a schema, we generate the proto description with comments for additional context for data generation. This component will also be able to extract column correlations from the SQL query body and annotated schema that would be useful in generating meaningful data.

### v3

# High-Fidelity and Complex Test Data Generation for Google SQL Code Generation Services

@@ Abstract
The demand for high-fidelity test data is paramount in industrial settings where access to production data is largely restricted. Traditional data generation methods often fall short, struggling with low-fidelity and the ability to model complex data structures and semantic relationships that are critical for testing complex SQL code generation services like Natural Language to SQL (NL2SQL). In this paper, we address the critical need for generating syntactically correct and semantically relevant high-fidelity mock data for complex data structures that includes columns with nested structures that we frequently encounter in Google workloads. We highlight the limitations of existing approaches used in production, particularly their inability to handle large and complex data structures, as well as the lack of semantically coherent test data that lead to limited test coverage. We demonstrate that by leveraging Large Language Models (LLMs) and incorporating strategic pre- and post-processing steps, we can generate syntactically correct and semantically relevant high-fidelity test data that adheres to complex structural constraints and maintains semantic integrity to the SQL test targets (queries/functions). This approach supports comprehensive testing of complex SQL queries involving joins, aggregations, and even deeply nested subqueries, ensuring robust evaluation of SQL code generation services, like NL2SQL and SQL Code Assistant. Our results demonstrate the practical utility of an LLM (gemini) based test data generation for industrial SQL code generation services where generating high-fidelity test data is essential due to the frequent unavailability and inaccessibility of production datasets for testing.
@@ I Introduction
One key challenge for effective test data generation in industrial contexts is ensuring high fidelity. High-fidelity data is syntactically correct and structurally compliant, accurately reflecting complex nested data structures and the nuanced relationships and correlations between fields and tables, as if it were sampled directly from production environments. This is crucial for supporting test cases with complex data structures that are consistent and satisfy data integrity requirements [5, 4, 9, 3]. For industrial applications heavily reliant on structured data types such as JSON, XML, protocol buffers (a.k.a. proto or protobuf [10]), it is paramount that the generated mock data strictly adheres to the specified structure and maintains data integrity and consistency [11, 12]. Specifically, our SQL code generation services utilize GoogleSQL [13] that supports complex nested data types, like array, struct and proto – making it a representative case for the challenges faced by many modern SQL dialects that support nested data structures. Conventional test data generation methods, such as manual test data generation by the developers, are highly susceptible to errors that can compromise the validity of testing and development. Furthermore, scalability and efficiency of generation are critical for enterprise needs, especially when a large amount of test data is needed [14, 15, 16]. As enterprise software systems grow in scale and complexity, the need for large volumes of mock data increases exponentially. Manual and heuristic data generation approaches become impractical and unsustainable in such environments, hindering development velocity and quality assurance efforts.
The generation also needs to ensure data value diversity to effectively simulate the wide range of scenarios encountered in real-world systems, including both common use cases and rare edge cases. Existing data generation methods often fall short in capturing this diversity, while rule-based systems tend to be inflexible and limited in their ability to generate varied data that reflects the complexities of industrial data distributions [17, 16].
@@
The emergence of Large Language Models (LLMs) presents a paradigm shift, offering new avenues for automating the generation of mock data, especially for structured formats such as database schema and protocol buffers. For clarity, throughout this paper, we use the term ”schema” to generally refer to a formal structural definition, encompassing both Google SQL schema definitions and/or Protocol Buffer (proto) definitions with potentially nested fields. LLMs can generate mock data that more closely approximates real-world data distributions and patterns based on their acquired real-world knowledge. LLMs can capture subtle nuances and semantic relationships between data attributes that are difficult to replicate with traditional methods [18, 19]. This is crucial for comprehensive production testing while also avoiding the privacy risks associated with using actual production data (regulatory security and privacy policy restricts access to user data). Additionally, LLMs can understand SQL structures and facilitate structurally compliant and schema-aware data generation [20, 21, 19, 22].
@@
While the use of LLMs for synthetic data generation is becoming popular in various domains [20, 14, 23, 24, 25, 26, 27, 28], the generation of structured data introduces unique challenges that are particularly relevant to industrial applications. A significant limitation of current LLM-based data generation approaches lies in their ability to effectively handle more complex data structures commonly found in our workloads. Specifically, LLMs struggle to accurately generate data that conforms to large (many columns or fields) and complex nested structures (Table II). While LLMs can generate data that adheres to basic and flat schema structures, accurately generating data that conforms to deeply nested structures (e.g., protobuf, GoogleSQL) with complex dependencies remains a challenge. This limitation is particularly problematic in testing real-world SQL generation services that rely on intricate data models, where the nuances of nested relationships over very wide tables are critical for comprehensive testing and validation in production.
@@
A further crucial challenge is ensuring the semantic validity of the generated data, especially when leveraging Large Language Models (LLMs). While LLMs might generate data that is syntactically and structurally correct, it often lacks the necessary semantic coherence across columns and tables. This occurs because many existing test data generation approaches focus primarily on approximating per-column or pairwise column distributions from production data [29, 30, 25, 17, 1, 3, 4, 5]. Whether using statistical or vanilla LLM methods [19, 22], this limited focus frequently leads to illogical combinations of column values, semantically incoherent join tables, and data inconsistencies—all of which violate real-world business rules.
Given a schema or proto definition, LLMs can be guided to generate mock data that adheres to the specified structure and constraints. However, tabular data generation via language models remains an active and evolving field, facing challenges such as accurately capturing complex inter-column (pairwise) relationships, mitigating the risk of “hallucinations” or data inconsistencies, and effectively handling rare values and outliers (also see Table II for vanilla LLM evaluation results) [19, 22]. Overcoming these limitations, including both structural and semantic challenges, is crucial for fully realizing the potential of LLMs in generating comprehensive and accurate mock data for complex workloads.
@@
In this paper, we address the critical need for generating syntactically correct and semantically relevant high-fidelity mock data for structurally well-defined, yet complex data formats, focusing on the specific requirements of industrial use cases. We define high-fidelity data as that which is syntactically accurate and also semantically aligned with the testing context, ensuring the generated data adheres to the constraints and logic implied by the test target. A key advantage is that, given a protocol buffer (proto) definition, LLMs can be guided to generate mock data that adheres to the specified structure and constraints. This includes respecting all implicit and explicit constraints, such as generating data within a specific date range when the SQL query under test includes a corresponding filter (e.g., WHERE year = 2022). This high-fidelity data is essential to support a wide range of complex test cases for services like NL2SQL and SQL code assistance systems. Robust evaluation of such generation systems requires this high-fidelity and complex test data, which is often inaccessible due to privacy concerns or a lack of sophisticated generation capabilities. Services like NL2SQL support arbitrarily complex use cases, generating intricate SQL queries involving joins, aggregations, and nested subqueries, for which vanilla out-of-the-box LLMs simply fail to generate suitable, high-quality test data.
@@
  • LLM-based High-Fidelity and Efficient Test Data Generation: We present a novel approach leveraging LLMs, enhanced with pre- and post-processing, to generate high-fidelity test data compliant with large complex structures, like production protobuf and GoogleSQL schema that include nested fields. Our generation method is significantly more efficient than manual generation used in production where production data access is restricted.
  • Enhanced Test Coverage for Complex SQL Queries: We demonstrate that our generated test data significantly improves test coverage of Google SQL code generation services, where many test cases are disabled due to limited access to production data and the difficulties of manual data generation. Our fundamental approach is SQL dialect-agnostic and should be readily applicable to other SQL code generation services that face similar challenges.
  • Semantic Coherence With Test SQL Queries: We emphasize and achieve the generation of semantically relevant test data. We provide a comparative analysis against production data sampling, showcasing the ability of our generated test data to cover edge cases and include semantically relevant predicate columns.
@@ II SQL code generation services
In this section, we examine the challenges of generating high-fidelity test data for two real-world GoogleSQL [13] SQL code generation services: SQL Code Assistant and NL2SQL. These services represent distinct use cases with varying complexities, highlighting the limitations of traditional test data generation methods which produces data that yields low coverage of the GoogleSQL queries under test.
TABLE I: Traditional test data generation methods for real-world SQL code generation services. Test coverage is largely limited by the test data quality, especially when the underlying data structure is large and complex with deep nesting.
@@
Table I summarizes test & schema complexities, as well as the traditional test data generation methods used for the services. The underlying data schema for both services are complex with many columns/fields with nested fields (nested proto and array types). The test complexity for NL2SQL is even more complex as the service generates SQL queries spanning over multiple tables across many customer domains. Test cases are submitted by the internal developers and also by crowdsourcing. The test SQL queries for SQL Code Assistant are mostly simple (single-table) selection queries; however, the manually curated test data overlooks complex nested structures from the schema limiting the test cases to SQL queries targeting mostly flat schemas with non-nested fields. The SQL query complexity vary, but most of them use nested fields and subqueries with joins. SQL code generation services require high-fidelity test data that adheres to the complex schema with nested structures that support complex and diverse sets of SQL queries/functions. Additionally, NL2SQL service evaluation requires testing if predicted SQL generates a result set that is semantically coherent with the natural language question. This also means that the test data instances should be semantically relevant and aligned with the context of the posed data question. Unfortunately, production sampling, though anonymized, is restricted to a few accessible domains and tables, resulting in incomplete test scenarios and low test coverage.
@@ II-A SQL Code Assistant
CREATE PUBLIC FUNCTION GetBalanceInUsd(Input fake_table) RETURNS INT64 AS ( CASE WHEN private_info.running_balance.currency = ’USD’ THEN private_info.running_balance.amount WHEN private_info.running_balance.currency = ’GBP’ THEN private_info.running_balance.amount * 1.26 WHEN private_info.running_balance.currency = ’EUR’ THEN private_info.running_balance.amount * 1.05 ELSE NULL END );
@@ II-B NL2SQL
//Question: Which fake_devices have tasks related to fake_task_name with a start date after 2020-06-01, and what are the tasks and their start dates for the fake_devices? ---------------------------------------------------- SELECT name, ARRAY_AGG(STRUCT(task, start_date)) FROM ( SELECT d.name, t.name AS task, FORMAT_TIMESTAMP("%Y-%m-%d", PARSE_TIMESTAMP("%Y-%m-%dT%H:%M:%E*S%Ez", t.start_date)) AS start_date FROM fake_project_tasks AS t INNER JOIN fake_devices AS d ON t.project_id = d.id WHERE ( t.name LIKE ’%fake_task_name’ ) AND t.start_date > "2020-06-01" ORDER BY start_date ) GROUP BY name;
@@ II-C Problem Definition
Let $\Lambda$ be a schema definition, like table schema or proto, with various types of fields/columns $\Xi$. The column set $\Xi$ includes primitive type (e.g., numerical, string, boolean) $\xi^{prim}$, categorical type $\xi^{cat}$ (e.g., enum), and nested type (e.g., struct, array, proto) $\xi^{nest}$ . $\xi^{nest}$ can be another (nested) schema definition $\Lambda_{i}$ that can have more nested structures within $\xi^{nest}_{i}$. Our work focuses on generating arbitrarily complex test data with any number of columns and with deeply nested structures, $\Xi=\{\xi^{prim}_{1},\xi^{cat}_{2},\xi^{nest}_{3},...,\xi_{n}\}$, for testing test SQL queries $\Omega=\{\omega_{1},\omega_{2},...,\omega_{m}\}$ and also without any training tabular dataset $\tau_{\Lambda}$ that confirms to $\Lambda$ – due to restricted production data access in real-world SQL code generation services.
@@ III-A Pre-processor
This module will seed the signals needed to generate high-fidelity data, automatically based on information from the context map and the test SQL query body. Additionally this can also be passed from the user. Any user provided input will override auto-generated ones. The signal will be a map of column name to column value (literal value, range or values, or a generator like incremental or other distribution, e.g., use a uniform distribution from 0 to 1 to populate this numeric column). These signals would be added to the context map and used later for post-processing.
Signals for Data Generation: Additionally the following signals would be needed to generate high-fidelity data:
@@
// Input SQL query: ---------------------------------------------------- SELECT * FROM ( SELECT * FROM fake_table_1 WHERE fake_column = "Regional_Team_Americas" AND date >= "2023-01-01" AND date <= "2023-03-31" ) AP LEFT JOIN (SELECT * FROM fake_table_2) AS Capacity ON SAFE_CAST(AP.date AS STRING) = Capacity.week_start_date AND AP.username = Capacity.username // Parsed/extracted structures: ---------------------------------------------------- Table Names: [] Column Names: ["*"] Constraints: $[$]] Subqueries: Table Names: ["fake_table_1"] Column Names: ["*"] Constraints: ["fake_column=’RegionalTeamAmericas’ANDdate>=’2023-01-01’ANDdate<=’2023-03-31’"] Table Names: ["fake_table_2"] Column Names: ["*"] Constraints: $[$]] // Join Constraints: ---------------------------------------------------- [{"fake_table_1": "date", "fake_table_2": "week_start_date"}, {"fake_table_1": "username", "fake_table_2": "username"}]
Figure 4: Pre-processor parses out sub-queries (middle) from the original SQL query (top) for seeding and independent and dependent data generation. Join information is extracted to highlight the table and column association for the join (bottom).
@@ III-B Context extractor
This module generates context relevant to the schema. We use a separate system to collect human written comments for a schema and proto from comments in data schema, source code and documentations. These comments are available at the schema/proto level and also individual column level. We use LLM to make these comments consistent and fill in missing columns. For protos, by themselves or as a column in a schema, we generate the proto description with comments for additional context for data generation. This component will also be able to extract column correlations from the SQL query body and annotated schema that would be useful in generating high-fidelity data.